Find parametric equations for line that is tangent to the curve x=cost, y=sint, z=t at the point
(cos(2π/6),sin(2π/6),2π/6) .

Parametrize the line so that it passes through the given point at t=0. All three answers are required for credit.
1
Expert's answer
2020-09-22T16:31:25-0400

"\displaystyle r(t) = (\cos{t}, \sin{t}, t)\\ 

\textsf{at}\hspace{0.1cm} t = 0, r(0) = (\cos{0}, \sin{0}, 0) = (1, 0, 0)\\

r'(t) = (-\sin{t}, \cos{t}, 1)\\


\textsf{At the point} \hspace{0.1cm}(x(t), y(t), z(t)) = \left(\cos\left(\frac{2\pi}{6}\right), \sin\left(\frac{2\pi}{6}\right), \frac{2\pi}{6}\right),\\

\textsf{ it corresponds to}\hspace{0.1cm} t = \frac{2\pi}{6}.\\

\textsf{Thus, we are looking for the}\\\textsf{tangent vector at}\hspace{0.1cm} t = \frac{2\pi}{6}\\

 
r'\left(\frac{2\pi}{6}\right)= \left(-\sin\left(\frac{2\pi}{6}\right), \cos\left(\frac{2\pi}{6}\right), 1\right)\\ = \left(-\frac{\sqrt{3}}{2},\frac{1}{2}, 1\right)\\

\therefore \textsf{The parametric equations for the line}\\\textsf{ that is tangent to the curve} \hspace{0.1cm} r(t) \hspace{0.1cm}\textsf{are}\\

x - 1 = -\frac{\sqrt{3}}{2}t, y - 0 = \frac{1}{2}t, z - 0 = t\\

x = -\frac{\sqrt{3}}{2}t + 1, y = \frac{1}{2}t, z = t"